OpenStudy (anonymous):

How many real zeroes does y = ( x + 8 )^3 + 9 have? (1 point) 1 2 3 0

OpenStudy (anonymous):

@hba

OpenStudy (campbell_st):

I think it only has 1 start with \[(x + 8)^3 + 9 = 0\] then \[(x + 8)^3 = -9\] and \[(x + 8) = \sqrt[3]{-9}\] should be easy to solve for x now...

OpenStudy (anonymous):

@campbell_st that didn't show up.

OpenStudy (campbell_st):

so did you solve for x or just look at the what I'd written..?

OpenStudy (anonymous):

Well, I couldn't see what you had written, it says [Math Processing Error], but I don't understand it.

OpenStudy (campbell_st):

ok.... well is you let y = 0 and solve for x you will find the number of zeros...